OpenStudy (anonymous):

Please Help Quick. Show how to solve seven-ninths times three-fourths with and without cancellation.

OpenStudy (anonymous):

\[\frac{7}{9} \times \frac{3}{4} = \frac{7 \times 3}{9 \times 4}\]